# Adversarial Robustness in Two-Stage Learning-to-Defer: Algorithms and Guarantees

### v1

@@ 5.2 Theoretical Guarantees
Theorem 10 establishes the consistency of the adversarial margin deferral surrogate family $\widetilde{\Phi}^{\rho,u}_{\text{def}}$, ensuring its alignment with the true adversarial deferral loss ~$\widetilde{\ell}_{\text{def}}$. The minimizability gaps derived in Theorem 10 vanish when $\mathcal{R}=\mathcal{R}_{\text{all}}$ and $\mathcal{G}=\mathcal{G}_{\text{all}}$ (Steinwart, 2007; Awasthi et al., 2022). Under the assumption that these gaps vanish, the following holds:
~$\mathcal{E}_{\widetilde{\ell}_{\text{def}}}(r,g)-\mathcal{E}_{\widetilde{\ell}_{\text{def}}}^{B}(\mathcal{R},\mathcal{G})\leq\mathcal{E}_{c_{0}}(g)-\mathcal{E}_{c_{0}}^{B}(\mathcal{G})+\Psi^{u}(1)\Big{(}\mathcal{E}_{\widetilde{\Phi}^{\rho,u}_{\text{def}}}(r)-\mathcal{E}_{\widetilde{\Phi}^{\rho,u}_{\text{def}}}^{\ast}(\mathcal{R})\Big{)}.$ (8)
After offline training of the multi-task model, we assume that the excess $c_{0}$-risk is bounded as $\mathcal{E}_{c_{0}}(g)-\mathcal{E}_{c_{0}}^{B}(\mathcal{G})\leq\epsilon_{0}$. Similarly, after training the rejector, the excess risk of the surrogate satisfies $\mathcal{E}_{\widetilde{\Phi}^{\rho,u}_{\text{def}}}(r)-\mathcal{E}_{\widetilde{\Phi}^{\rho,u}_{\text{def}}}^{\ast}(\mathcal{R})\leq\epsilon_{1}$. Under these conditions, the $\ell_{\text{def}}$-excess risk is bounded as ~$\mathcal{E}_{\widetilde{\ell}_{\text{def}}}(r,g)-\mathcal{E}_{\widetilde{\ell}_{\text{def}}}^{B}(\mathcal{R},\mathcal{G})\leq\epsilon_{0}+\Psi^{u}(1)\epsilon_{1}$, establishing both the Bayes-consistency and $(\mathcal{R},\mathcal{G})$-consistency of the surrogate losses $\widetilde{\Phi}^{\rho,u}_{\text{def}}$.
Building on the theoretical guarantees of the non-convex family $\widetilde{\Phi}^{\rho,u}_{\text{def}}$, we introduce a smooth adversarial regularized deferral (SARD) algorithm. SARD extends the standard SAD framework by incorporating a regularization term that enhances stability and robustness. Despite this modification, SARD preserves the key theoretical guarantees of $\widetilde{\Phi}^{\rho,u}_{\text{def}}$, ensuring consistency and minimizability under the same conditions (Mao et al., 2023b; Awasthi et al., 2023).
Guarantees for SARD:
Using the fact that $\widetilde{\Phi}^{\text{smth},u}_{01}\geq\widetilde{\Phi}^{\rho,u,j}_{01}$, we establish guarantees for our smooth adversarial deferral surrogate family $\widetilde{\Phi}^{\text{smth},u}_{\text{def}}$ under the same conditions.
Corollary 11 (Guarantees for SAD).
Assume $\mathcal{R}$ is symmetric and locally $\rho$-consistent. Then, for the agent set $\mathcal{A}$, any hypothesis $r\in\mathcal{R}$, and any distribution $\mathcal{D}$, the following holds for a multi-task model $g\in\mathcal{G}$:
$\quad+\Psi^{u}(1)\Big{(}\mathcal{E}_{\widetilde{\Phi}^{\text{smth},u}_{\text{def}}}(r)-\mathcal{E}_{\widetilde{\Phi}^{\rho,u}_{\text{def}}}^{\ast}(\mathcal{R})+\mathcal{U}_{\widetilde{\Phi}^{\rho,u}_{\text{def}}}(\mathcal{R})\Big{)}$
@@
Corollary 11 establishes that SAD shows similar consistency properties under the given conditions, with the minimizability gap vanishing for $\mathcal{R}=\mathcal{R}_{\text{all}}$ and $\mathcal{G}=\mathcal{G}_{\text{all}}$. This motivates the development of an adversarial robustness algorithm based on minimizing a regularized empirical formulation of SAD, referred to as SARD.
Proposition 12 (SARD: Smooth Adversarial Regularized Deferral Algorithm).
Assume $\mathcal{R}$ is symmetric and locally $\rho$-consistent. For a regularizer $\Omega$ and hyperparameter $\eta>0$, the regularized empirical risk minimization problem for SARD is:
$\min_{r\in\mathcal{R}}\Bigg{[}\frac{1}{K}\sum_{k=1}^{K}\widetilde{\Phi}^{\text{smth},u}_{\text{def}}(r,g,m,z_{k})+\eta\Omega(r)\Bigg{]}.$
@@ 6.1 Multiclass Classification Task
Categories were assigned to three experts with a correctness probability $p=0.94$, while the remaining probability was uniformly distributed across the other categories, following the approach in (Mozannar and Sontag, 2020; Verma et al., 2022; Cao et al., 2024). To further evaluate robustness, we introduced a weak expert M3, with only a few assigned categories, and assumed that the attacker is aware of this weakness. Agent costs are defined as $c_{0}(h(x),y)=\ell_{01}(h(x),y)$ for the model and $c_{j>0}(m_{j}^{h}(x),y)=\ell_{01}(m_{j}^{h}(x),y)$, aligned with (Mozannar and Sontag, 2020; Mozannar et al., 2023; Verma et al., 2022; Cao et al., 2024; Mao et al., 2023a). Both the model and the rejector were implemented using ResNet-4 (He et al., 2015). The agents’ performance, additional training details, and experimental results are provided in Appendix E.1.
@@ 6.2 Regression Task
We evaluate the performance of SARD against the method proposed by Mao et al. (2024d) using the California Housing dataset involving median house price prediction (Kelley Pace and Barry, 1997).
@@
We train three experts, each implemented as an MLP, specializing in a specific subset of the dataset based on a predefined localization criterion. Among these, expert M3 is designed to specialize in a smaller region, resulting in comparatively weaker overall performance. Agent costs for regression are defined as $c_{0}(f(x),t)=\text{RMSE}(f(x),t)$ for the model and $c_{j>0}(m_{j}^{f}(x),t)=\text{RMSE}(m_{j}^{f}(x),t)$, aligned with (Mao et al., 2024d). Both the model and the rejector are trained on the full dataset using MLPs. We provide detailed agent performance results, training procedures, and additional experimental details in Appendix E.2.
@@
Table 2: Performance comparison of SARD with the baseline (Mao et al., 2024d) on the California Housing dataset. The table reports Root Mean Square Error (RMSE).
@@ 6.3 Multi-Task
We train two Faster R-CNN models (Ren et al., 2016) as experts, each specializing in a distinct subset of the dataset. Expert M1 is trained exclusively on images containing animals, while expert M2 focuses on images with vehicles. Agent costs are defined as $c_{0}(g(x),z)=\text{mAP}(g(x),z)$ for the model and $c_{j>0}(m_{j}(x),z)=\text{mAP}(m_{j}(x),z)$, aligned with (Montreuil et al., 2024). The primary model and the rejector are implemented as lightweight versions of Faster R-CNN using MobileNet (Howard et al., 2017). We provide detailed performance results, training procedures, and additional experimental details in Appendix E.3.
@@ Appendix A Notation and Preliminaries for the Appendices
  • $\psi:\mathcal{Y}\times\mathcal{T}\times\mathcal{Z}\to\mathbb{R}^{+}$: Quantify the prediction’s quality.
@@
Appendix B Attacks Illustration
B.1 Untargeted Attack
@@ Appendix C Algorithm
  Return: trained rejector model $r^{\ast}$.
@@ Appendix D Proof Adversarial Robustness in Two-Stage Learning-to-Defer
D.1 Proof Lemma 4
See 4
Proof  In adversarial training, the objective is to optimize the worst-case scenario of the objective function under adversarial inputs $x^{\prime}\in B_{p}(x,\gamma)$. For our case, we start with the standard L2D loss for the two-stage setting (Mao et al., 2024d; Montreuil et al., 2024):
@@ D.1 Proof Lemma 4
D.2 Proof Lemma 5
See 5
Proof  Referring to adversarial true deferral loss defined in Lemma 4, we have:
@@ D.2 Proof Lemma 5
D.3 Proof Lemma 6
See 6
Proof  Let $x\in\mathcal{X}$ denote an input and $x_{j}^{\prime}\in B_{p}(x,\gamma)$ an adversarially perturbed input within an $\ell_{p}$-norm ball of radius $\gamma$ for each agent. Let $r\in\mathcal{R}$ be a rejector. We now define the composite-sum $\rho$-margin losses for both clean and adversarial scenarios:
@@ D.3 Proof Lemma 6
D.4 Proof Lemma 7
See 7
Proof  Using Lemma 5, we have:
@@ D.4 Proof Lemma 7
Since $\widetilde{\Phi}^{\rho,u,j}_{01}\leq\widetilde{\Phi}^{\text{smth},u}_{01}$ by Lemma 6, we obtain:
@@
D.5 Proof Lemma 9
See 9
Proof
@@ D.5 Proof Lemma 9
$=\sum_{i=0}^{J}\sup_{\xi_{i}^{\prime}\in B_{p}(x,\gamma)}p(x,\{i\}_{\xi_{i}^{\prime}}^{r^{\ast}})\Psi^{u}\Bigg{(}\sum_{j=0}^{i-1}\Psi_{\rho}\Big{(}r^{\ast}(\xi_{i}^{\prime},\{i\}^{r^{\ast}}_{\xi_{i}^{\prime}})-r^{\ast}(\xi^{\prime}_{i},\{j\}^{r^{\ast}}_{\xi_{i}^{\prime}})\Big{)}+J-i\Bigg{)}\quad\text{($\Psi_{\rho}(t)=1,\forall t\leq 0$)}$
$=\sum_{i=0}^{J}\sup_{\xi_{i}^{\prime}\in B_{p}(x,\gamma)}p(x,\{i\}_{\xi_{i}^{\prime}}^{r^{\ast}})\Psi^{u}(J-i)\quad\text{($\Psi_{\rho}(v)=0,\forall v\geq\rho$ and $\inf_{x^{\prime}_{j}\in B_{p}(x,\gamma)}|r^{*}(x^{\prime}_{j},q)-r^{*}(x^{\prime}_{j},q^{\prime})|\geq\rho$)}$
$=\sum_{i=0}^{J}p_{[i]}(x)\Psi^{u}(J-i)\quad\text{($r^{\ast}$ and $p(x)$ same ordering of $j\in\mathcal{A}$)}$
@@
$\geq\sum_{i=0}^{J}\sup_{\xi_{i}^{\prime}\in B_{p}(x,\gamma)}p(x,\{i\}_{\xi_{i}^{\prime}}^{r})\Psi^{u}\Bigg{(}\sum_{j=0}^{i-1}\Psi_{\rho}\Big{(}r(\xi_{i}^{\prime},\{i\}_{\xi_{i}^{\prime}})-r(\xi^{\prime}_{i},\{j\}_{\xi_{i}^{\prime}})\Big{)}+J-i\Bigg{)}-\Big{(}\sum_{i=0}^{J}p_{[i]}(x)\Psi^{u}(J-i)\Big{)}$
@@
$\geq\Psi^{u}(1)p(x,r(x))1_{r\notin\overline{\mathcal{R}}_{\gamma}(x)}-\sum_{i=0}^{J}p_{[i]}(x)\Psi^{u}(J-i)+\sum_{i=0}^{J}p(x,\{i\}^{r}_{x})\Psi^{u}(J-i)\quad(\text{sup}_{\xi_{i}^{\prime}\in B_{p}(x,\gamma)}p(x,\{i\}^{r}_{\xi_{i}^{\prime}})\geq p(x,\{i\}^{r}_{x}))$
$=\Psi^{u}(1)p(x,r(x))1_{r\notin\overline{\mathcal{R}}_{\gamma}(x)}+\Psi^{u}(1)\Bigg{(}\max_{j\in\mathcal{A}}p(x,j)-p(x,r(x))\Bigg{)}+\begin{bmatrix}\Psi^{u}(1)\\ \Psi^{u}(1)\\ \Psi^{u}(2)\\ \vdots\\ \Psi^{u}(J)\end{bmatrix}\cdot\begin{bmatrix}p(x,\{J\}^{r}_{x})\\ p(x,\{J-1\}^{r}_{x})\\ p(x,\{J-2\}^{r}_{x})\\ \vdots\\ p(x,\{0\}^{r}_{x})\end{bmatrix}$
$\quad-\begin{bmatrix}\Psi^{u}(1)\\ \Psi^{u}(1)\\ \Psi^{u}(2)\\ \vdots\\ \Psi^{u}(J)\end{bmatrix}\cdot\begin{bmatrix}p_{[J]}(x)\\ p_{[J-1]}(x)\\ p_{[J-2]}(x)\\ \vdots\\ p_{[0]}(x)\end{bmatrix}$
@@
D.6 Proof Theorem 10
See 10
Proof  Using Lemma 5, we have:
@@ D.6 Proof Theorem 10
Then, using Lemma 9, it leads to:

### v2

@@ 5.2 Theoretical Guarantees
Theorem 5.10 establishes that the adversarial margin deferral surrogate family $\widetilde{\Phi}^{\rho,u}_{\text{def}}$ is consistent with the adversarial true deferral loss ~$\widetilde{\ell}_{\text{def}}$. When the minimizability gaps vanish—i.e., when $\mathcal{R}=\mathcal{R}_{\text{all}}$ and $\mathcal{G}=\mathcal{G}_{\text{all}}$ (Steinwart, 2007; Awasthi et al., 2022)—the bound reduces to:
~$\mathcal{E}_{\widetilde{\ell}_{\text{def}}}(r,g)-\mathcal{E}_{\widetilde{\ell}_{\text{def}}}^{B}(\mathcal{R},\mathcal{G})$ $\leq\mathcal{E}_{c_{0}}(g)-\mathcal{E}_{c_{0}}^{B}(\mathcal{G})$
$+\Psi^{u}(1)\left(\mathcal{E}_{\widetilde{\Phi}^{\rho,u}_{\text{def}}}(r)-\mathcal{E}_{\widetilde{\Phi}^{\rho,u}_{\text{def}}}^{\ast}(\mathcal{R})\right).$
Assume that after offline training, the multi-task model satisfies $\mathcal{E}_{c_{0}}(g)-\mathcal{E}_{c_{0}}^{B}(\mathcal{G})\leq\epsilon_{0}$, and the trained rejector satisfies $\mathcal{E}_{\widetilde{\Phi}^{\rho,u}_{\text{def}}}(r)-\mathcal{E}_{\widetilde{\Phi}^{\rho,u}_{\text{def}}}^{\ast}(\mathcal{R})\leq\epsilon_{1}$. Then, the excess adversarial deferral risk is bounded as:
~$\mathcal{E}_{\widetilde{\ell}_{\text{def}}}(r,g)-\mathcal{E}_{\widetilde{\ell}_{\text{def}}}^{B}(\mathcal{R},\mathcal{G})\leq\epsilon_{0}+\Psi^{u}(1)\epsilon_{1},$
thereby establishing both Bayes-consistency and $(\mathcal{R},\mathcal{G})$-consistency of the surrogate losses $\widetilde{\Phi}^{\rho,u}_{\text{def}}$.
We now introduce the smooth adversarial regularized deferral (SARD) algorithm, which extends the standard SAD framework by incorporating a regularization term. SARD retains the theoretical guarantees of $\widetilde{\Phi}^{\rho,u}_{\text{def}}$, including consistency and minimizability under the same assumptions (Mao et al., 2023b; Awasthi et al., 2023).
Guarantees for SARD.
We now extend our theoretical guarantees to the smooth adversarial deferral surrogate (SAD) family $\widetilde{\Phi}^{\text{smth},u}_{\text{def}}$, leveraging the inequality $\widetilde{\Phi}^{\text{smth},u}_{01}\geq\widetilde{\Phi}^{\rho,u,j}_{01}$. Under the same assumptions as in Theorem 5.10, we obtain the following bound:
Corollary 5.11 (Guarantees for SAD).
Assume $\mathcal{R}$ is symmetric and locally $\rho$-consistent. Then, for the agent set $\mathcal{A}$, any hypothesis $r\in\mathcal{R}$, and any distribution $\mathcal{D}$, the following holds for any multi-task model $g\in\mathcal{G}$:
~$\mathcal{E}_{\widetilde{\ell}_{\text{def}}}(r,g)-\mathcal{E}_{\widetilde{\ell}_{\text{def}}}^{B}(\mathcal{R},\mathcal{G})+\mathcal{U}_{\widetilde{\ell}_{\text{def}}}(\mathcal{R},\mathcal{G})$
$\leq\Psi^{u}(1)\left(\mathcal{E}_{\widetilde{\Phi}^{\text{smth},u}_{\text{def}}}(r)-\mathcal{E}_{\widetilde{\Phi}^{\rho,u}_{\text{def}}}^{\ast}(\mathcal{R})+\mathcal{U}_{\widetilde{\Phi}^{\rho,u}_{\text{def}}}(\mathcal{R})\right)$
@@
Corollary 5.11 shows that SAD satisfies similar consistency guarantees under the same assumptions as its non-smooth counterpart. In particular, the minimizability gap vanishes when $\mathcal{R}=\mathcal{R}_{\text{all}}$ and $\mathcal{G}=\mathcal{G}_{\text{all}}$, further justifying the use of SAD in practice.
Motivated by this theoretical foundation, we introduce a learning algorithm that minimizes a regularized empirical approximation of SAD, recovering guarantees established in both Theorem 5.10 and Corollary 5.11. We refer to this algorithm as SARD (Smooth Adversarial Regularized Deferral).
Proposition 5.12 (SARD: Smooth Adversarial Regularized Deferral Algorithm).
Assume $\mathcal{R}$ is symmetric and locally $\rho$-consistent. Let $\Omega:\mathcal{R}\to\mathbb{R}^{+}$ be a regularizer and $\eta>0$ a hyperparameter. The regularized empirical risk minimization objective solved by SARD is:
$\min_{r\in\mathcal{R}}\left[\frac{1}{K}\sum_{k=1}^{K}\widetilde{\Phi}^{\text{smth},u}_{\text{def}}(r,g,m,z_{k})+\eta\,\Omega(r)\right].$
@@
5.3 Computational complexity of SARD
Having established statistical guarantees, we now quantify the complexity cost of SARD. Because training time can become a bottleneck for large‑scale models, especially when inner maximizations are involved, we spell out the precise forward/backward counts so that future work can reproduce or improve our implementation.
Proposition 5.13 (Epoch cost of SARD).
[…]
$n\,\bigl{(}1+|\mathcal{A}|\,T_{a}\bigr{)}\bigl{(}C_{\mathrm{fwd}}+C_{\mathrm{bwd}}\bigr{)}$
network traversals, while the peak memory equals that of a single forward–backward pass plus the storage of one adversarial copy of each input.
Proof.
Per mini‑batch SARD executes (i) one clean forward of $r_{\theta}$; (ii) for each $j\in\mathcal{A}$ and each of the $T_{a}$ PGD steps, one forward and one backward of $r_{\theta}$; (iii) one backward pass to update $\theta$. This totals $2(1+|\mathcal{A}|\,T_{a})$ network traversals. Multiplying by the $\lceil n/B\rceil$ batches yields (5.13). Because PGD steps are processed sequentially, only the activations of the clean forward plus those of the current PGD iteration must reside in memory, giving the stated peak footprint. ∎
Comparison with Two-Stage L2D losses.
For the standard convex deferral loss $\Phi_{\mathrm{def}}^{u}=\sum_{j\in\mathcal{A}}\tau_{j}\,\Phi_{01}^{u}\bigl{(}r,x,j\bigr{)}$ the inner maximizations is absent ($T_{a}=0$), so its epoch cost is simply $n\,(C_{\mathrm{fwd}}+C_{\mathrm{bwd}})$. Equation (5.13) therefore shows that SARD introduces a purely multiplicative overhead factor $1+|\mathcal{A}|\,T_{a}$.
Although the factor in (5.13) appears large, the absolute overhead is modest on contemporary hardware. Moreover, the overhead scales linearly with both the number of experts $|\mathcal{A}|$ and the attack depth $T_{a}$, providing clear knobs for future work to trade robustness for wall‑time. These observations guide our experimental choices in the next section.
@@ 6.1 Multiclass Classification Task
Categories were assigned to three experts with a correctness probability $p=0.94$, while the remaining probability was uniformly distributed across the other categories, following the approach in (Mozannar & Sontag, 2020; Verma et al., 2022; Cao et al., 2024). To further evaluate robustness, we introduced a weak expert M3, with only a few assigned categories, and assumed that the attacker is aware of this weakness. Agent costs are defined as $c_{0}(h(x),y)=\ell_{01}(h(x),y)$ for the model and $c_{j>0}(m_{j}^{h}(x),y)=\ell_{01}(m_{j}^{h}(x),y)$, aligned with (Mozannar & Sontag, 2020; Mozannar et al., 2023; Verma et al., 2022; Cao et al., 2024; Mao et al., 2023a). Both the model and the rejector were implemented using ResNet-4 (He et al., 2015). The agents’ performance, additional training details, and experimental results are provided in Appendix E.1.
@@ 6.2 Regression Task
We evaluate the performance of SARD against the method proposed by Mao et al. (2024c) using the California Housing dataset involving median house price prediction (Kelley Pace & Barry, 1997).
@@
We train three experts, each implemented as an MLP, specializing in a specific subset of the dataset based on a predefined localization criterion. Among these, expert M3 is designed to specialize in a smaller region, resulting in comparatively weaker overall performance. Agent costs for regression are defined as $c_{0}(f(x),t)=\text{RMSE}(f(x),t)$ for the model and $c_{j>0}(m_{j}^{f}(x),t)=\text{RMSE}(m_{j}^{f}(x),t)$, aligned with (Mao et al., 2024c). Both the model and the rejector are trained on the full dataset using MLPs. We provide detailed agent performance results, training procedures, and additional experimental details in Appendix E.2.
@@
Table 2: Performance comparison of SARD with the baseline (Mao et al., 2024c) on the California Housing dataset. The table reports Root Mean Square Error (RMSE).
@@ 6.3 Multi-Task
We train two Faster R-CNN models (Ren et al., 2015) as experts, each specializing in a distinct subset of the dataset. Expert M1 is trained exclusively on images containing animals, while expert M2 focuses on images with vehicles. Agent costs are defined as $c_{0}(g(x),z)=\text{mAP}(g(x),z)$ for the model and $c_{j>0}(m_{j}(x),z)=\text{mAP}(m_{j}(x),z)$, aligned with (Montreuil et al., 2024). The primary model and the rejector are implemented as lightweight versions of Faster R-CNN using MobileNet (Howard et al., 2017). We provide detailed performance results, training procedures, and additional experimental details in Appendix E.3.
@@ 7 Conclusion
Acknowledgment
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD-2023-01-041-J) and by A*STAR, and is part of the programme DesCartes which is supported by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme.
Impact Statement
This paper introduces methods to improve the adversarial robustness of two-stage Learning-to-Defer frameworks, which allocate decision-making tasks between AI systems and human experts. The work has the potential to advance the field of Machine Learning, particularly in high-stakes domains such as healthcare, finance, and safety-critical systems, where robustness and reliability are essential. By mitigating vulnerabilities to adversarial attacks, this research ensures more secure and trustworthy decision-making processes.
The societal implications of this work are largely positive, as it contributes to enhancing the reliability and fairness of AI systems. However, as with any advancement in adversarial robustness, there is a potential for misuse if adversarial strategies are exploited for harmful purposes. While this paper does not directly address these ethical concerns, we encourage further exploration of safeguards and responsible deployment practices in future research.
No immediate or significant ethical risks have been identified in this work, and its societal impacts align with the well-established benefits of improving robustness in Machine Learning systems.
@@ Appendix A Notation and Preliminaries for the Appendices
  • $\psi:\mathcal{Y}\times\mathcal{T}\times\mathcal{Y}\times\mathcal{T}\to\mathbb{R}^{+}$: Quantify the prediction’s quality.
@@
Appendix B Illustration
B.1 Learning-to-Defer in Inference
Figure 1: Learning-to-Defer in inference. The input $x$ is routed to the most cost-effective agent in the system by the rejector $r\in\mathcal{R}$: in this case, the human expert.
B.2 Untargeted Attack
@@ Appendix C Algorithm
  Return: trained rejector model $\hat{r}$.
@@ Appendix D Proof Adversarial Robustness in Two-Stage Learning-to-Defer
D.1 Proof Lemma 5.2
See 5.2
Proof.
In adversarial training, the objective is to optimize the worst-case scenario of the objective function under adversarial inputs $x^{\prime}\in B_{p}(x,\gamma)$. For our case, we start with the standard L2D loss for the two-stage setting (Mao et al., 2024c; Montreuil et al., 2024):
@@
D.2 Proof Lemma 5.5
See 5.5
Proof.
Referring to adversarial true deferral loss defined in Lemma 5.2, we have:
@@
D.3 Proof Lemma 5.6
See 5.6
Proof.
Let $x\in\mathcal{X}$ denote an input and $x_{j}^{\prime}\in B_{p}(x,\gamma)$ an adversarially perturbed input within an $\ell_{p}$-norm ball of radius $\gamma$ for each agent. Let $r\in\mathcal{R}$ be a rejector. We now define the composite-sum $\rho$-margin losses for both clean and adversarial scenarios:
@@
D.4 Proof Lemma 5.7
See 5.7
Proof.
Using Lemma 5.5, we have:
@@
Since $\widetilde{\Phi}^{\rho,u,j}_{01}\leq\widetilde{\Phi}^{\text{smth},u}_{01}$ by Lemma 5.6, we obtain:
@@
D.5 Proof Lemma 5.9
See 5.9
Proof.
@@
$=\sum_{i=0}^{J}\sup_{\xi_{i}^{\prime}\in B_{p}(x,\gamma)}p(x,\{i\}_{\xi_{i}^{\prime}}^{r^{\ast}})\Psi^{u}\Bigg{(}\sum_{j=0}^{i-1}\Psi_{\rho}\Big{(}r^{\ast}(\xi_{i}^{\prime},\{i\}^{r^{\ast}}_{\xi_{i}^{\prime}})-r^{\ast}(\xi^{\prime}_{i},\{j\}^{r^{\ast}}_{\xi_{i}^{\prime}})\Big{)}+J-i\Bigg{)}\quad\text{$\Big{(}\Psi_{\rho}(t)=1,\forall t\leq 0\Big{)}$}$
$=\sum_{i=0}^{J}\sup_{\xi_{i}^{\prime}\in B_{p}(x,\gamma)}p(x,\{i\}_{\xi_{i}^{\prime}}^{r^{\ast}})\Psi^{u}(J-i)\quad\text{$\Big{(}\Psi_{\rho}(v)=0,\forall v\geq\rho$ and $\inf_{x^{\prime}_{j}\in B_{p}(x,\gamma)}|r^{*}(x^{\prime}_{j},q)-r^{*}(x^{\prime}_{j},q^{\prime})|\geq\rho\Big{)}$}$
$=\sum_{i=0}^{J}p_{[i]}(x)\Psi^{u}(J-i)\quad\text{$\Big{(}r^{\ast}$ and $p(x)$ same ordering of $j\in\mathcal{A}\Big{)}$}$
@@
$\geq\sum_{i=0}^{J}\sup_{\xi_{i}^{\prime}\in B_{p}(x,\gamma)}p(x,\{i\}_{\xi_{i}^{\prime}}^{r})\Psi^{u}\Bigg{(}\sum_{j=0}^{i-1}\Psi_{\rho}\Big{(}r(\xi_{i}^{\prime},\{i\}_{\xi_{i}^{\prime}})-r(\xi^{\prime}_{i},\{j\}_{\xi_{i}^{\prime}})\Big{)}+J-i\Bigg{)}$
$-\Big{(}\sum_{i=0}^{J}p_{[i]}(x)\Psi^{u}(J-i)\Big{)}$
@@
$\geq\Psi^{u}(1)p(x,r(x))1_{r\notin\overline{\mathcal{R}}_{\gamma}(x)}-\sum_{i=0}^{J}p_{[i]}(x)\Psi^{u}(J-i)+\sum_{i=0}^{J}p(x,\{i\}^{r}_{x})\Psi^{u}(J-i)$
$\quad(\text{sup}_{\xi_{i}^{\prime}\in B_{p}(x,\gamma)}p(x,\{i\}^{r}_{\xi_{i}^{\prime}})\geq p(x,\{i\}^{r}_{x})$
Then RHS is equal to the above,
$\Delta\mathcal{C}_{\widetilde{\Phi}_{01}^{\rho,u,j}}(r,x)$ $\geq\Psi^{u}(1)p(x,r(x))1_{r\notin\overline{\mathcal{R}}_{\gamma}(x)}+\Psi^{u}(1)\Bigg{(}\max_{j\in\mathcal{A}}p(x,j)-p(x,r(x))\Bigg{)}+\begin{bmatrix}\Psi^{u}(1)\\ \Psi^{u}(1)\\ \Psi^{u}(2)\\ \vdots\\ \Psi^{u}(J)\end{bmatrix}\cdot\begin{bmatrix}p(x,\{J\}^{r}_{x})\\ p(x,\{J-1\}^{r}_{x})\\ p(x,\{J-2\}^{r}_{x})\\ \vdots\\ p(x,\{0\}^{r}_{x})\end{bmatrix}$ (23)
$\quad-\begin{bmatrix}\Psi^{u}(1)\\ \Psi^{u}(1)\\ \Psi^{u}(2)\\ \vdots\\ \Psi^{u}(J)\end{bmatrix}\cdot\begin{bmatrix}p_{[J]}(x)\\ p_{[J-1]}(x)\\ p_{[J-2]}(x)\\ \vdots\\ p_{[0]}(x)\end{bmatrix}$ (24)
@@
D.6 Proof Theorem 5.10
See 5.10
Proof.
Using Lemma 5.5, we have:
@@
Then, using Lemma 5.9, it leads to: